### v1

@@ 5.2 Multiple Superimposed Oscillators : known dynamics
The task was originally introduced in a presentation by Jaeger, 2004b , for $n=2$ (what would now be called MSO-2, but there called “additive dynamics”), as an example of a task that an ESN finds “impossible to learn”. Wierstra et al., (2005) increase task difficulty by extending the list of $\alpha$ with higher frequency values to give MSO-5, and Roeschies and Igel, 2010b further extend the $\alpha$ values to give MSO-8. Other authors who used this benchmark include Xue et al., (2007) (MSO-2) and Koryakin et al., (2012) (MSO-8).
Since Jaeger, 2004b states that the task cannot be solved by the standard ESN model, some authors use this benchmark on extensions to the ESN model (Wierstra et al.,, 2005; Schmidhuber et al.,, 2007; Xue et al.,, 2007; Koryakin et al.,, 2012), for example, by evolving the weights, or using feedback, to demonstrate superior performance.
However, if one uses just the standard ESN model with similar training and testing dataset lengths as these variants (many hundreds of timesteps each), one can also achieve apparently reasonable NRMSE values (see section 9.3.1), implying that the ESN has learned the data. Jaeger et al., 2007a (, fig.4) demonstrate that this good prediction does not hold indefinitely; after many thousands of timesteps, the prediction diverges. This is what Jaeger, 2004b means when referring to the task as “impossible to learn”.
@@ 5.3 Lazy Figure 8 : known dynamics
The figure 8 task is a “perennial task” in the field of Recurrent Neural Networks (Jaeger et al., 2007a, ), and a challenging one. It involves predicting the next point along a sequence that traces the shape of the digit 8, and as such is a prediction of a known system.
In Reservoir Computing, this task is adapted to the “Lazy [^7] While the “Lazy” in “Lazy Eight” typically refers to the orientation of the figure 8 drawn on its side ($\infty$), Jaeger instead calls it thus due to the timescale over which the figure is drawn. Figure 8” (Jaeger et al., 2007a, ), where the full shape is composed of 200 points. Training is performed over 3000 steps, which includes a washout period of 1000 steps. Two experiments were performed, the first where the 200 points were equally spaced over the figure, and a second “time-warped” where the spacing of the points varied over time.
Like the MSO task, this benchmark is one that cannot be satisfactorily performed by a classical ESN, although Jaeger et al., 2007a show that their leaky–integrator based model is capable of accomplishing it. This is an example of a benchmark being used to show that a newer model performs better than classical ESNs.
Figure 8: Lazy Figure 8 (Jaeger et al., 2007a, , fig.6). This figure follows a trajectory of 200 points, which is then repeated over a training and subsequent testing period. The two figures are of the output from the ESNs, with (left) constant and (right) “time-warped” inputs.
@@ 5.4.1 Lorenz’63 system
The first system, introduced in 1963 (Lorenz,, 1963, eqn.25–27), is frequently referred to as Lorenz’63. The set of three coupled ODEs are:
@@
The values of the parameters $\sigma$, $\rho$, and $\beta$ change behaviour between periodic and chaotic. The parameter values used are typically $\sigma=10$, $\rho=28$, $\beta=8/3$ (Pathak et al.,, 2017; Roeschies and Igel, 2010a, ; Canaday et al.,, 2021). The behaviour for this chaotic choice is illustrated in figure 9(a).
@@ 5.4.2 Lorenz’96 system
The second system, introduced in 1996 (Lorenz,, 1995, eqn.1), is typically referred to as Lorenz’96. It is described by $N$ variables $x_{i}$, $i\in 0\ldots N-1$, governed by $N$ coupled ODEs:
@@ 5.5 Kuramoto–Sivashinski equation: known dynamics
The Kuramoto–Sivashinsky (KS) equation is a fourth order PDE originally derived to model diffusion-induced chemical turbulence (Sivashinsky,, 1977, 1980). When used as a benchmark, the task is to predict the next output of the system, and therefore is a prediction task for a known system.
@@
where $y_{t}=\partial y/\partial t$ and $y_{x}=\partial y/\partial x$. The initial condition (at $t=0$) is usually defined on a finite $x$ domain ($0\leq x<L)$ with periodic boundary conditions. The time behaviour is periodic for small $L$; as $L$ is increased, the periodic behaviour undergoes bifurcations, and eventually becoming chaotic.
The KS equation has been used in the context of Reservoir Computing. Initially it was used as an actual task (Pathak et al.,, 2017): here the Reservoir Computer’s task is to predict the next output of the equation, to test if the Reservoir Computer can imitate the Lyapunov exponents of the equation, whether or not divergence takes place. The K–S equation is used to find out whether Reservoir Computing could solve the observer problem, which involves deducing the full state of a dynamical system based on partial information of the state and of the system dynamics (Lu et al.,, 2017).
The KS equation has also been used as a prediction benchmark. Vlachas et al., (2020) uses it with a spatial size of $L=200$; Rafayelyan et al., (2020) use $L\in[12,22,36,60,100]$.
Pathak et al., (2018) also compare against a modified form, to investigate the effect of spatial inhomogeneity:
@@
The size $L$ is an integer multiple of the wavelength $\lambda$, with $L\in[100,200,400,800,1600]$.
@@ 5.6 Sunspot numbers : unknown dynamics
Sunspots are dark spots that appear on the sun on a temporary basis. Sunspots have been recorded consistently since 1610 (NGDC,, 2014). As such, the observations can be used as a dataset for the prediction of the behaviour of an unknown dynamical system.
Predicting the next value in a dataset of sunspots is a perennial task in Machine Learning, first used in the early twentieth century (Yule,, 1927). The task varies greatly in the details: while each involves predicting the next value in the dataset, which dataset is used, how the values are calculated, and what preprocessing is applied, all vary across the literature. The only consistency appears to be that the datasets typically start in 1749, although there remains some inconsistency on whether the first value should be taken in January (Stepney,, 2021) or July (Schwenker and Labib,, 2009; Rodan and Tino,, 2011).
@@
NOAA dataset NGDC, (2014) Schwenker and Labib, (2009)
@@
SILSO, (2022) Shougat et al., (2021)
[…]
Table 5: Sources found in the literature for the sunspot numbers, and an example citation of the source’s use in the literature as the “Sunspot Prediction” benchmark.
@@
At least four different sources for “Sunspot Numbers” are used in the literature, as shown in table 5; the Zurich dataset is shown in figure 11. These have differing ways of dealing with missing data. In the NGDC dataset, a value of $-99$ indicated missing data; other datasets appear to have removed this data in preprocessing. Some authors who use the NGDC dataset refer to preprocessing to deal with missing data, but do not define how this is performed (Schwenker and Labib,, 2009).
As well as the existence of differing datasets, other problems with this task are found. There is the inconsistency of the time period used: some use months (Stepney,, 2021), others days (Shougat et al.,, 2021). There does not seem to be any specification on how monthly averages are computed, and whether or how they account for the different lengths of months. There is also the impossibility of making perfect observations: issues may arise from cloud cover leading to missing data, quality of telescopes, or difficulty deciding on how to count sunspots. All these render this data unreliable as data about a dynamical system, and make comparison between different systems problematic.
That said, finding out whether a data-driven model driven by imperfect data can produce usable results is a potentially valuable problem; all data is, after all, imperfect: data is merely a representation of the real state, and like all representation, some details are abstracted away. Having a dataset where some of the limitations are known and acknowledged thus has its own value; but if this is the case, it is important for users of the dataset to acknowledge this.
@@ 5.7 Santa Fe laser dataset: unknown dynamics
The Santa Fe Lasers dataset is a dataset originally produced in an experiment trying to replicate Lorenz-like chaos in NH3–FIR LASERs (Hübner et al.,, 1994). It was initially distributed as one of the datasets in the Santa Fe 1992 Time Series prediction competition, the proceedings of which were published in 1994 (Weigend and Gershenfeld,, 1994). The dataset is now frequently used as a prediction task for an unknown system.
@@
The most readily available version of the dataset, with the original 1000 values, can be found at Salles, (2021)888the full dataset is available on github: https://github.com/MaterialMan/CHARC/blob/master/Support%20files/other/Datasets/laser.txt. The amount of data used varies in the literature, and is occasionally not recorded, as we can see in the examples listed in table 6.
@@
Larger et al., (2012) not recorded not recorded not recorded
Brunner et al., (2013) 4000 3200 800
Dale, 2018b 5000 25000
@@
Guo et al., (2019) 4000 3000 1000
Shougat et al., (2021) not recorded not recorded not recorded
[…]
@@
The Santa Fe 1992 Time Series Prediction competition also distributed five other datasets. These are not used as benchmarks, except for the sleep apnea dataset (see section 7.4). It is generally safe to assume that any work referring to the “Santa Fe benchmark” with no further elaboration is using the laser dataset.
@@ 6 Computation tasks
6.1 the XOR Task : binary output
The XOR task is a family of tasks based on the bitwise operation XOR, and entered the Reservoir Computing field through the field of Neural Networks. Unlike many of the benchmarks discussed here, there is no single XOR task, nor consensus on what the goal of the task is. Here we define it as a single-output classification task, and describe various variants, to demonstrate the ways the task can be used.
An XOR task is used early on in Reservoir Computing, where a bucket of water is shown to support a simple Liquid State Machines (Fernando and Sojakka,, 2003). The task to output the XOR of the two simultaneously applied inputs. Much post-processing is needed to read the output. The task of outputting the XOR of two consecutively supplied input bits is also used (Laporte et al.,, 2021; Shougat et al.,, 2021).
The XOR task has been adapted to measure the memory of a Reservoir Computer as well as the nonlinearity (Verstraeten et al.,, 2010): the task is to output the XOR of two initial values after a sizeable delay. A similar task has been used to benchmark the memory of an ESN (Jaeger,, 2012).
@@ 6.2 Parity task : binary output
Given a stream of inputs $u(t)$, the output of the PARITY-$n$ task is given by the parity of $n$ consecutive bits, read out after a delay $\tau$ (Haykin et al.,, 2007):
$y(t)=PARITY(u(t-\tau),u(t-\tau-1),...u(t-\tau-n))$ (19)
The input $u(t)$ is randomly sampled from $[-1,1]$ (Dion et al.,, 2018) or $[0,1]$(Haykin et al.,, 2007).
@@ 7.1 MNIST handwritten digits : static classification
The MNIST Handwritten Digits Database (LeCun et al.,, 1998) is a database of digits between 0 and 9 handwritten by highschool students and United States Census Bureau employees. The database consists of multiple datasets, most notable among which is the training dataset, composed of 60,000 images created by 250 writers. While originally black and white, the images were normalised to fit into $20$ by $20$ pixel arrays, and subsequently centered in a $28$ by $28$ pixel image, with the resulting images being in grayscale.
This task has been used as a Reservoir Computing benchmark. (Du et al.,, 2017) describe a preprocessing technique that converts the image of the digit into a black-and-white $22\times 20$ pixel image. Each row is then converted into a time series input fed into the reservoir. In order to make the task easier for the reservoir, it takes two inputs, sampled at different rates. It is unclear here how the data is fed in to the reservoir.
Another use of this database when benchmarking Reservoir Computing involves turning each image into a time series (Schaetti et al.,, 2016), in this case feeding the image to the reservoir a single pixel at a time. Another variation is to permute the pixels, in order to remove some of the internal structures of the image. This variation is called the Permuted Sequential MNIST task (psMNIST) (Manneschi et al., 2021b, ; Manneschi et al., 2021a, ).
@@ 7.2 Spoken digits : non-stationary classification
The TI-46 Isolated Spoken Words Dataset (Doddington and Schalk,, 1981) is a dataset of 20 individual words, spoken by 8 men and 8 women. It contains ten digits, and ten words common in speech recognition, such as “help” and “stop”. The dataset as available has 26 utterances per female speaker of each word digit, 10 for training and 16 for testing. The intent of the project was to have a dataset of words that personal assistant-style technology could be tested on, particularly ones suited to office environments, rather than home ones. The dataset can be found online (Liberman et al.,, 1980), but is not open access.
A subset of the TI-46 Dataset consisting of the digits from “zero” to “nine” said by five different speakers was first used in the context of Reservoir Computing to the speech recognition capabilities of Liquid State Machines to the then state-of-the-art Hidden Markov Models (Verstraeten et al.,, 2005). This work was expanded to include ESNs (Verstraeten et al.,, 2006). The authors detail the experimental setup, which we summarise here:
  • Noise was added to the words, using the NOISEX database of Varga and Steeneken, (1993).
  • The words were first preprocessed using the Lyon Passive Ear Model (Lyon,, 1982), a more biologically inspired model for preprocessing than the one performed for Hidden Markov Models. (Some other authors preprocess the data differently.)
  • The words were then encoded into spike trains using the Bens Spiker Algorithm (BSA) (Schrauwen and Van Campenhout,, 2003)
@@
Verstraeten et al., (2006) added a degree of confidence measure to the output, which increased the accuracy of their results.
The Isolated Spoken Digits task is by far the most popular classification benchmark, particularly after its first use. See, for example, Verstraeten et al., (2007); Butcher et al., (2010); Appeltant et al., (2011); Duport et al., 2012a ; Larger et al., (2012); Paquot et al., (2012); Brunner et al., (2013); Dejonkheere et al., (2014); Soriano et al., (2015); Vinckier et al., (2015); Dion et al., (2018); Moon et al., (2019).
A simple version of this benchmark was used in one of the earliest reservoir computing publications (Fernando and Sojakka,, 2003), getting a Reservoir Computer to recognise the spoken digits “0” and “1”, with these digits being input as recorded speech samples.
Isolated spoken digits used as a standard benchmark, as described here, is separate from the larger problem of general speech recognition, which attempts to identify words and phonemes from larger datasets (Triefenbach et al.,, 2010).

### v2

@@ 5.2 Multiple Superimposed Oscillators: known dynamics
The task was originally introduced in a presentation by Jaeger [73], for $n=2$ (what would now be called MSO-2, but there called ‘additive dynamics’), as an example of a task that an ESN finds ‘impossible to learn’. The task difficulty has been increased by extending the list of $\alpha$ with higher frequency values to give MSO-5 [74] and MSO-8 [75]. Other uses of this benchmark include [76] (MSO-2) and [77] (MSO-8).
Since Jaeger [73] states that the task cannot be solved by the standard ESN model, some authors use this benchmark on extensions to the ESN model [74, 78, 76, 77], for example, by evolving the weights, or using feedback, to demonstrate superior performance. However, if one uses just the standard ESN model with similar training and testing dataset lengths as these variants (many hundreds of timesteps each), one can also achieve apparently reasonable NRMSE values (see section 9.3.1), implying that the ESN has learned the data. Jaeger et al. [12, fig.4] demonstrate that this good prediction does not hold indefinitely: after many thousands of timesteps, the prediction diverges. This is what Jaeger [73] means when referring to the task as ‘impossible to learn’.
@@ 5.3 Lazy Figure 8: known dynamics
The figure 8 task is a ‘perennial task’ in the field of Recurrent Neural Networks [12], and a challenging one. It involves predicting the next point along a sequence that traces the shape of the digit 8, and as such is a prediction of a known system.
In Reservoir Computing, this task is adapted to the ‘Lazy figure 8’ [^7] While the ‘Lazy’ in ‘Lazy Eight’ typically refers to the orientation of the figure 8 drawn on its side ($\infty$), Jaeger instead calls it thus due to the timescale over which the figure is drawn. [12], where the full shape is composed of 200 points. Training is performed over 3000 steps, which includes a washout period of 1000 steps. Two experiments were performed, the first where the 200 points were equally spaced over the figure, and a second ‘time-warped’ where the spacing of the points varied over time.
Like the MSO task, this benchmark is one that cannot be satisfactorily performed by a classical ESN, although Jaeger et al. [12] show that their leaky integrator based model is capable of accomplishing it. This is an example of a benchmark being used to show that a newer model performs better than classical ESNs.
Figure 8: Lazy Figure 8 [12, fig.6]. This figure follows a trajectory of 200 points, which is then repeated over a training and subsequent testing period. The two figures are of the output from the ESNs, with (left) constant and (right) ‘time-warped’ inputs.
@@ 5.4.1 Lorenz’63 system
The first system, introduced in 1963 [79, eqn.25–27], is frequently referred to as Lorenz’63. The set of three coupled ODEs are:
@@
The values of the parameters $\sigma$, $\rho$, and $\beta$ change behaviour between periodic and chaotic. The parameter values used are typically $\sigma=10$, $\rho=28$, $\beta=8/3$ [80, 71, 81]. The behaviour for this chaotic choice is illustrated in figure 9(a).
@@ 5.4.2 Lorenz’96 system
The second system, introduced in 1996 [82, eqn.1], is typically referred to as Lorenz’96. It is described by $N$ variables $x_{i}$, $i\in 0\ldots N-1$, governed by $N$ coupled ODEs:
@@ 5.5 Kuramoto–Sivashinski equation: known dynamics
The Kuramoto–Sivashinsky (KS) equation is a fourth order PDE originally derived to model diffusion-induced chemical turbulence [85, 86]. When used as a benchmark, the task is to predict the next output of the system, and therefore is a prediction task for a known system.
@@
where $y_{t}=\partial y/\partial t$ and $y_{x}=\partial y/\partial x$. The initial condition ($t=0$) is usually defined on a finite $x$ domain ($0\leq x<L)$ with periodic boundary conditions. The time behaviour is periodic for small $L$; as $L$ is increased, the periodic behaviour undergoes bifurcations, and eventually becoming chaotic (Figure 10).
The KS equation has been used in the context of Reservoir Computing. Initially it was used as an actual task [80]: here the Reservoir Computer’s task is to predict the next output of the equation, to test if the Reservoir Computer can imitate the Lyapunov exponents of the equation, whether or not divergence takes place. The KS equation is used to find out whether Reservoir Computing could solve the observer problem, which involves deducing the full state of a dynamical system based on partial information of the state and of the system dynamics [87].
The KS equation has also been used as a prediction benchmark, with spatial sizes of $L=200$ [84] and $L\in\{12,22,36,60,100\}$ [88].
Pathak et al. [89] also compare against a modified form, to investigate the effect of spatial inhomogeneity:
@@
where the size $L$ is an integer multiple of the wavelength $\lambda$, with $L\in\{100,200,400,800,1600\}$.
@@ 5.6 Sunspot numbers: unknown dynamics
Sunspots are dark spots that appear on the sun on a temporary basis. Sunspots have been recorded consistently since 1610 [90]. As such, the observations can be used as a dataset for the prediction of the behaviour of an unknown dynamical system.
Predicting the next value in a dataset of sunspots is a perennial task in Machine Learning, first used in the early twentieth century [91]. The task varies greatly in the details: while each involves predicting the next value in the dataset, which dataset is used, how the values are calculated, and what preprocessing is applied, all vary across the literature. The only consistency appears to be that the datasets typically start in 1749, although there remains some inconsistency on whether the first value should be taken in January [21] or July [92, 37].
@@
NOAA dataset [90] [92]
@@
[95] [26]
Table 5: Sources found in the literature for the sunspot numbers, and an example citation of the source’s use in the literature as the ‘Sunspot Prediction’ benchmark.
@@
At least four different sources for sunspot numbers are used in the literature, as shown in table 5; the Zurich dataset is shown in figure 11. These have differing ways of dealing with missing data. In the NGDC dataset, a value of $-99$ indicated missing data; other datasets appear to have removed this data in preprocessing. Some authors who use the NGDC dataset refer to preprocessing to deal with missing data, but do not define how this is performed [92].
As well as the existence of differing datasets, other problems with this task are found. There is the inconsistency of the time period used: some use months [21], others days [26]. There does not seem to be any specification on how monthly averages are computed, and whether or how they account for the different lengths of months. There is also the impossibility of making perfect observations: issues may arise from cloud cover leading to missing data, quality of telescopes, or difficulty deciding on how to count sunspots. All these render this data unreliable as data about a dynamical system, and make comparison between different systems problematic.
That said, finding out whether a data-driven model driven by imperfect data can produce usable results is a potentially valuable problem; all real world data is, after all, imperfect: data is merely a representation of the real state, and like all representation, some details are abstracted away. Having a dataset where some of the limitations are known and acknowledged thus has its own value; but if this is the case, it is important for users of the dataset to acknowledge this.
@@ 5.7 Santa Fe Laser dataset: unknown dynamics
The Santa Fe Laser dataset is a dataset originally produced in an experiment trying to replicate Lorenz-like chaos in NH3–FIR lasers [96]. It was initially distributed as one of the datasets in the Santa Fe 1992 Time Series prediction competition, the proceedings of which were published in 1994 [32]. The dataset is now frequently used as a prediction task for an unknown system.
@@
The most readily available version of the dataset, with the original 1000 values, can be found at [97]888the full dataset is available from: https://github.com/MaterialMan/CHARC/blob/master/Support%20files/other/Datasets/laser.txt . The amount of data used varies in the literature, and is occasionally not recorded, as we can see in the examples listed in table 6.
@@
[98] not recorded not recorded not recorded
[99] 4000 3200 800
[45] 5000 25000
@@
[100] 4000 3000 1000
[26] not recorded not recorded not recorded
@@
The Santa Fe 1992 Time Series Prediction competition also distributed five other datasets. These are not used as benchmarks, except for the sleep apnea dataset (see section 7.4). It is generally safe to assume that any work referring to the ‘Santa Fe benchmark’ with no further elaboration is using the laser dataset.
5.8 McMaster IPIX radar data : unknown dynamics
The McMaster IPIX Radar is an X-band radar, originally created to detect icebergs, but capable of collecting data on all sea clutter. The website [101] indicates that the movement of sea clutter is influenced by several dynamical systems, and is itself a nonlinear dynamical system. As such, it makes an ideal prediction benchmark for Reservoir Computing.
The radar’s website makes two databases available for general use [101], one created from data recorded in Dartmouth, Nova Scotia, in 1993, and the other from data collected in Grimsby, Ontario, in 1998. The latter database is marked as incomplete by the radar’s website, and is encouraged to be used only comparatively with the Dartmouth data.
This data has been used as a prediction task in Reservoir Computing [76]. It is unclear exactly which of the datasets are used, or how these are provided as input for the Reservoir Computer. 2000 datapoints are used, of which 200 are used as washout data, 800 for training, and 1000 for testing, but no more information is given.
Subsequent uses of this dataset in Reservoir Computing [102, 14, 37, 51, 50] appear to be consistent with the first use, citing [76] and using the same dataset sizes.
@@ 6 Computation tasks
6.1 XOR Task: binary output
The XOR task is a family of tasks based on the bitwise operation XOR, and entered the Reservoir Computing field through the field of Neural Networks. Unlike many of the benchmarks discussed here, there is no single XOR task, nor consensus on what the goal of the task is. Here we define it as a single-output classification task, and describe some variants, to demonstrate the ways the task can be used.
An XOR task is used early in the history of Reservoir Computing, where a bucket of water is shown to support a simple Liquid State Machine [16]. The task is to output the XOR of the two simultaneously applied inputs. Much post-processing is needed to read the output. The task of outputting the XOR of two consecutively supplied input bits is also used [103, 26].
The XOR task has been adapted to measure the memory of a Reservoir Computer as well as the nonlinearity [104]: the task is to output the XOR of two initial values after a sizeable delay. A similar task has been used to benchmark the memory of an ESN [7].
@@ 6.2 Parity task: binary output
Given a stream of inputs $u(t)$, the output of the PARITY-$n$ task is given by the parity of $n$ consecutive bits, read out after a delay $\tau$ [105]:
[…]
The input $u(t)$ is randomly sampled from $[-1,1]$ [106] or $[0,1]$[105].
@@ 7.1 MNIST handwritten digits : static classification
The MNIST Handwritten Digits Database [108] is a database of digits between 0 and 9 handwritten by high-school students and United States Census Bureau employees. The database consists of multiple datasets, most notable among which is the training dataset, composed of 60,000 images created by 250 writers. While originally black and white, the images were normalised to fit into $20\times 20$ pixel arrays, and subsequently centred in a $28\times 28$ pixel image, with the resulting images being in greyscale.
This task has been used as a Reservoir Computing benchmark. [109] describe a preprocessing technique that converts the image of the digit into a black-and-white $22\times 20$ pixel image. Each row is then converted into a time series input fed into the reservoir. In order to make the task easier for the reservoir, it takes two inputs, sampled at different rates. It is unclear here how the data is fed in to the reservoir.
Another use of this database when benchmarking Reservoir Computing involves turning each image into a time series [110], in this case feeding the image to the reservoir a single pixel at a time. Another variation is to permute the pixels, in order to remove some of the internal structures of the image. This variation is called the Permuted Sequential MNIST task (psMNIST) [111, 112].
@@ 7.2 Spoken digits: non-stationary classification
The TI-46 Isolated Spoken Words Dataset [27] is a dataset of 20 individual words, spoken by 8 men and 8 women. It contains ten digits, and ten words common in speech recognition, such as ‘help’ and ‘stop’. The dataset as available has 26 utterances per female speaker of each word digit, 10 for training and 16 for testing. The intent of the project was to have a dataset of words that personal assistant-style technology could be tested on, particularly ones suited to office environments, rather than home ones. The dataset can be found online [113], but is not open access.
A subset of the TI-46 Dataset consisting of the digits from ‘zero’ to ‘nine’ said by five different speakers was first used in the context of Reservoir Computing to the speech recognition capabilities of Liquid State Machines to the then state-of-the-art Hidden Markov Models [6]. This work was expanded to include ESNs [114]. The authors detail the experimental setup, which we summarise here:
  • Noise was added to the words, using the NOISEX database [115].
  • The words were first preprocessed using the Lyon Passive Ear Model [116], a more biologically inspired model for preprocessing than the one performed for Hidden Markov Models. (Some other authors preprocess the data differently.)
  • The words were then encoded into spike trains using the Bens Spiker Algorithm (BSA) [117].
@@
Verstraeten et al. [114] added a degree of confidence measure to the output, which increased the accuracy of their results.
The Isolated Spoken Digits task is by far the most popular classification benchmark, particularly after its first use. See, for example, [38, 118, 13, 51, 98, 40, 99, 50, 119, 42, 106, 72].
A simple version of this benchmark was used in one of the earliest reservoir computing publications [16], getting a Reservoir Computer to recognise the spoken digits “0” and “1”, with these digits being input as recorded speech samples.
Isolated spoken digits used as a standard benchmark, as described here, is separate from the larger problem of general speech recognition, which attempts to identify words and phonemes from larger datasets [10].